Joe​ Henry's machine shop uses 2,520 brackets during the course of a year. These brackets are purchased from a supplier 90 miles away. The following information is known about the​ brackets:

Annual demand

2,520

Holding cost per bracket per year

$1.25

Order cost per order

$18.50

Lead time

22 days

Working days per year

250

a)

What is the​ EOQ? ____units ​(round your response to two decimal​ places).

​b)

What is the average inventory if the EOQ is​ used? _____units ​(round your response to two decimal​ places).
What would be the annual inventory holding​ cost? $_____(round your response to two decimal​ places).

​c)

Given the​ EOQ, how many orders will be made​ annually? ____orders ​(round your response to two decimal​ places).

What would be the annual order​ cost? $_____​(round your response to two decimal​ places).

​d)

Given the​ EOQ, what is the total annual cost of managing​ (ordering and​ holding) the​ inventory? $____(round your response to two decimal​ places).

​e) What is the time between​ orders? _____days ​(round your response to two decimal​ places).

f)

What is the reorder point​ (ROP)? ____units ​(round your response to two decimal​ places).
Answer #1

a) EOQ = Sqroot{(2*Annual Demand* Ordering cost per order)/Holding cost per unit}

= Sqroot{(2*2520*18.5)/1.25}

= 273.12

b) Average Inventory = EOQ/2

= 273.12/2

= 136.56

c) Number of orders annualy = Annual Demand/EOQ

= (2520/273.12)

= 9.23

d) Annual Cost of Inventory = Annual Ordering cost + Annual Holding cost

Annual ordering cost = Number of orders * Ordering cost per order

= (9.23*18.5)

= $170.76

Annual Holding cost = Average Inventory * Holding cost per unit

= (136.56*1.25)

= $170.7

Annual cost of inventory = $170.76 + $170.7

= $341.46

e) Time between orders = Number of working days/Number of orders

= (250/9.23)

= 27.09 days

f) Reorder Point = Average daily demand * Delivery lead time

= (2520/250)*22

= $221.76 units

Please provide up votes, if you are satisfied with the response. Comment in case of any issue regarding the same.